If i dilute 75mL of .41M CaF2 solution to a final volume of 375, what will be the molarity of the resulting solution?
1
Expert's answer
2020-05-07T14:16:00-0400

1) "\nu" (for the 1st solution) = "C\u2022V" = 41(M) • 0.075(L) = 3.08 mol ; the amount of substance from dilution doesn't change "\to" "\nu(1)=\nu(2)" = 3.08 mol

2)"C" (for solution after dilution) = "\nu/V" = 3.08(mol) / 0.375 (L) = 8.21 M

Answer: C (for the solution after dilution) = 8.21 M